# EventFlow: Forecasting Temporal Point Processes with Flow Matching

### v2

@@ Appendix B Proofs
Let $A_{1},A_{2}\subseteq\Gamma$ be Borel measurable (Daley & Vere-Jones, 2003, Prop. 5.3) subsets of the configuration space $\Gamma$, i.e. each of $A_{1},A_{2}$ is a measurable collection of event sequences. Observe that for $i=1,2$, each $A_{i}$ can be written as a disjoint union
@@ Appendix C EventFlow Architecture and Training Details
Here, we provide additional details regarding the parametrization and training of our EventFlow model. In general, our model is based on the transformer architecture (Vaswani et al., 2017; Yang et al., 2022), due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions. We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture, i.e., other choices like CNNs or RNNs are certainly possible here. In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$. All models (including the baselines) were trained on a cluster of six NVIDIA A6000 GPUs with 24Gb of VRAM. No hardware parallelization was used, i.e., each model was only trained on a single GPU.
@@
Our unconditional model takes in a sequence $\gamma_{s}=(t^{1}_{s},\dots,t^{n}_{s})$ and flow time $s$. We first embed the sequence times $t^{k}$, the flow-time $s$, and the sequence position indices. These position indices are handled by sinusoidal embeddings followed by an additional linear layer. There are three linear layers in total – one for the flow time, one shared across the sequence times, and one for the position indices. These embeddings are added together to create a representation for each element of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$. Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$. See Figure 3.
For the conditional model, we use a standard transformer encoder-decoder architecture. We first embed the history sequence times $\mathcal{H}$ and the sequence position indices in a manner analogous to the above. In addition, the model was provided the start of the prediction window $T_{0}$ by concatenating it as the final event in $\mathcal{H}$. This yielded better results than encoding the start of the prediction window separately. We feed these embeddings through the transformer encoder produce an intermediate representation $e_{\mathcal{H}}$.
@@
Our architecture for predicting the number of future events given a history, i.e. $p_{\phi}(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model. However, the key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit. See Figure 5. We note that this requires specifying a maximum value $N_{\max}$, which we set as the maximum sequence length seen in the training data. In early experiments, we also parametrized the model not through a logit but rather as a mixture of Poissons or a mixture of Negative Binomials. While this no longer has the limitation of assuming a value for $N_{\max}$, we found that this approach yielded worse results.
@@
We normalize all sequences to the range $[-1,1]$, using the overall min/max event time seen in the training data. All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation. Our models are trained with the Adam (Kingma & Ba, 2015) optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler (Loshchilov & Hutter, 2017), which cycled every $10,000$ steps. Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
To tune our model, we performed a grid search over learning rates in $\{5\times 10^{-3},10^{-3},5\times 10^{-4}\}$ and dropout probabilities in $\{0,0.1,0.2\}$. Overall, we found that learning rates of $10^{-2}$ or larger often caused the model to diverge, and a dropout of $0.1$ yielded the best results across all settings. We use $6$ transformer layers, $8$ attention heads, and an embedding dimension of $512$ across all settings, except for the Reddit-C and Reddit-S datasets where we use $4$ heads and an embedding dimension of $128$ due to the increased memory cost of these datasets.
@@
Learning Rate Emb. Dim. MLP Dim Heads Transformer Layers
Hawkes1 $10^{-3}$ $512$ $2048$ $8$ $6$
Hawkes2 $10^{-3}$ $512$ $2048$ $8$ $6$
Nonstationary Poisson $10^{-3}$ $512$ $2048$ $8$ $6$
Nonstationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
Stationary Renewal $10^{-3}$ $512$ $2048$ $8$ $6$
Self-Correcting $10^{-3}$ $512$ $2048$ $8$ $6$
PUBG $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $5\times 10^{-3}$ $128$ $256$ $4$ $6$
Taxi $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Twitter $10^{-3}$ $512$ $2048$ $8$ $6$
Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
@@
Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $10^{-3}$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $128$ $256$ $4$ $6$
Taxi $10^{-3}$ $512$ $2048$ $8$ $6$
Twitter $5\times 10^{-4}$ $512$ $2048$ $8$ $6$
Yelp-Airport $10^{-3}$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $512$ $2048$ $8$ $6$
@@
Learning Rate $\alpha\times N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers
PUBG $10^{-3}$ $1.0$ $512$ $2048$ $8$ $6$
Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ $4$ $6$
Taxi $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Twitter $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ $8$ $6$
Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ $8$ $6$
@@ Appendix D Additional Details on Baselines
In this section, we provide additional details regarding our baseline methods. All methods are trained at a batch size of $64$ for $1,000$ epochs, using early stopping on the validation set loss. In early experiments, we also evaluated AttNHP (Zuo et al., 2020), a variant of the NHP which uses an attention-based encoder, but found it to be prohibitively expensive in terms of memory cost (requiring more than 24 GB of VRAM) and, as a result, do not include it in our results.
@@
Our first baseline is the intensity-free TPP model of Shchur et al. (2020a). This model uses an RNN encoder and a mixture of log-normal distributions to parametrize the decoder. We directly use the implementation provided by the authors. [^2] URL: https://github.com/shchur/ifl-tpp. We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and mixture component counts $\{8,16,32,64\}$. Our best hyperparameters can be found in Table 8 and Table 9.
@@
Hawkes1 $10^{-3}$ $10^{-4}$ $32$ $8$
Hawkes2 $10^{-2}$ 00 $32$ $8$
Nonstationary Poisson $10^{-3}$ $10^{-6}$ $128$ $8$
@@
Reddit-C $10^{-2}$ 00 $64$ $8$
Reddit-S $10^{-2}$ 00 $64$ $16$
Taxi $10^{-3}$ $10^{-6}$ $128$ $8$
Twitter $10^{-2}$ $10^{-5}$ $32$ $8$
@@
We additionally compare against the Neural Hawkes Process of Mei & Eisner (2017). This model uses an LSTM encoder and a parametric form, whose weights are modeled by a neural network, to model the conditional intensity function. In practice, we use the implementation proved by the EasyTPP benchmark (Xue et al., 2024), as this version implements the necessary thinning algorithm for sampling. [^3] URL: https://github.com/ant-research/EasyTemporalPointProcess We perform a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and embedding dimensions in $\{32,64,128\}$. These hyperparameters are chosen as the EasyTPP implementation allows these to be configured easily. Our best hyperparameters are reported in Table 10 and Table 11.
@@
Our diffusion baseline is based on the implementation of Lin et al. (2022), and our decoder model architecture is taken directly from the code of Lin et al. (2022). [^4] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model (Ho et al., 2020) trained to generate a single inter-arrival time given a history embedding. Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (2) is not tractable. Instead, the model is trained by maximizing an ELBO of only the subsequent inter-arrival time.
In preliminary experiments, we found that the codebase provided by Lin et al. (2022) often produced NaN values during sampling, prompting us to make several changes. First, we use the RNN encoder from Shchur et al. (2020a), i.e. the same encoder as the IFTPP baseline, to reduce the memory requirements of the model. Second, we do not log-scale the inter-arrival times as suggested by Lin et al. (2022), as we found that this often led to overflow and underflow issues at sampling time. Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range $[-1,1]$. This is aligned with standard diffusion implementations (Ho et al., 2020), and allows us to perform clipping at sampling time to avoid the accumulation of errors. With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported (Lüdke et al., 2023).
We use $1000$ diffusion steps and the cosine beta schedule (Nichol & Dhariwal, 2021), and we train the model on the simplified $\epsilon$-prediction loss of Ho et al. (2020). We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and layer numbers $\{2,4,6\}$. Our best hyperparameters can be found in Table 12 and Table 13.
@@
Hawkes1 $10^{-3}$ $10^{-6}$ $64$ $2$
Hawkes2 $10^{-2}$ $10^{-5}$ $64$ $4$
Nonstationary Poisson $10^{-3}$ $10^{-5}$ $128$ $2$
Nonstationary Renewal $10^{-3}$ $10^{-4}$ $64$ $2$
Stationary Renewal $10^{-2}$ 00 $32$ $6$
Self-Correcting $10^{-3}$ 00 $32$ $6$
PUBG $10^{-3}$ 00 $64$ $2$
Reddit-C $10^{-3}$ $10^{-6}$ $128$ $4$
Reddit-S $10^{-3}$ 00 $64$ $4$
Taxi $10^{-2}$ 00 $128$ $4$
Twitter $10^{-3}$ $10^{-4}$ $64$ $6$
Yelp-Airport $10^{-2}$ 00 $32$ $2$
Yelp-Miss. $10^{-2}$ $10^{-5}$ $128$ $2$
@@
PUBG $10^{-4}$ $10^{-5}$ $32$ $6$
Reddit-C $10^{-2}$ $10^{-6}$ $64$ $6$
Reddit-S $10^{-3}$ 00 $64$ $4$
Taxi $10^{-3}$ $10^{-6}$ $32$ $2$
Twitter $10^{-4}$ $10^{-5}$ $64$ $6$
Yelp-Airport $10^{-4}$ $10^{-5}$ $64$ $6$
Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ $4$
@@
We compare to the Add-and-Thin model of Lüdke et al. (2023) as a recently proposed non-autoregressive baseline. We directly run the code provided by the authors without additional modifications. [^5] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al. (2023), in order to ensure a fair comparison between the methods considered. We train for $1,000$ epochs with early stopping on the validation loss. Tuning is performed via a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and number of mixture components in $\{8,16,32,64\}$. We choose to tune only these hyperparameters in order to follow the implementation provided by the authors. Our best hyperparameters can be found in Table 14 and Table 15.
@@
Nonstationary Renewal $10^{-2}$ $8$
Stationary Renewal $10^{-2}$ $8$
Self-Correcting $10^{-4}$ $8$
PUBG $10^{-3}$ $8$

### v3

@@ Appendix B Proofs
Let $A_{1},A_{2}\subseteq\Gamma$ be Borel measurable [Daley and Vere-Jones, 2003, Prop. 5.3] subsets of the configuration space $\Gamma$, i.e. each of $A_{1},A_{2}$ is a measurable collection of event sequences. Observe that for $i=1,2$, each $A_{i}$ can be written as a disjoint union
@@ Appendix C EventFlow Architecture and Training Details
Here, we provide additional details regarding the parametrization and training of our EventFlow model. Our model is based on the transformer architecture [Vaswani et al., 2017, Yang et al., 2022], due to its general ability to handle variable length inputs and outputs, high flexibility, and ability to incorporate long-range interactions. We emphasize that this is an effective architecture choice, but our method is not necessarily tied to this architecture. In all settings, our reference measure $\mu_{0}$ is specified with $q=\mathcal{N}(0,I)$. All models (including the baselines) were trained on a cluster of six NVIDIA A6000 GPUs with 24Gb of VRAM. No hardware parallelization was used, i.e., each model was only trained on a single GPU.
@@
Our unconditional model takes in a sequence $\gamma_{s}=(t^{1}_{s},\dots,t^{n}_{s})$ and flow time $s$. We first embed the sequence times $t^{k}$, the flow-time $s$, and the sequence position indices. These position indices are handled by sinusoidal embeddings followed by an additional linear layer. There are three linear layers in total: one for the flow time, one shared across the sequence times, and one for the position indices. These embeddings are added together to create a representation for each element of the sequence, and we apply a standard transformer to this sequence to produce a sequence of vectors of length $N(\gamma_{s})$. Finally, each of these vectors is projected to one dimension via a final linear layer with shared weights to produce the vector field $v_{\theta}(\gamma_{s},s)$. See Figure 3.
For the conditional model, we use a standard transformer encoder-decoder architecture. We first embed the history sequence times $\mathcal{H}$ and the sequence position indices in a manner analogous to the above. The model was provided the start of the prediction window $T_{0}$ by concatenating it as the final event in $\mathcal{H}$. This yielded better results than encoding the start of the prediction window separately. We feed these embeddings through the transformer encoder produce an intermediate representation $e_{\mathcal{H}}$.
@@
Our architecture for predicting the number of future events given a history, i.e. $p_{\phi}(n\mid\mathcal{H})$, is again based on the transformer decoder, sharing the same overall architecture as our unconditional model. The key difference is that we instead take a mean of the final sequence embeddings before passing this through a small MLP to produce the final logit. See Figure 5. We note that this requires specifying a maximum value $N_{\max}$, which we set as the maximum sequence length seen in the training data. In early experiments, we also parametrized the model not through a logit but rather as a mixture of Poissons or a mixture of Negative Binomials. While this no longer has the limitation of assuming a value for $N_{\max}$, we found that this approach yielded worse results.
@@
We normalize all sequences to the range [1,1][-1,1], using the overall min/max event time seen in the training data. All sequences are generated on this normalized scale, prior to re-scaling the sequence back to the original data range before evaluation. Our models are trained with the Adam [Kingma and Ba, 2015] optimizer with $\beta_{1}=0.9$ and $\beta_{2}=0.999$ for $30,000$ steps with a cosine scheduler [Loshchilov and Hutter, 2017], which cycled every $10,000$ steps. Final hyperparameters were selected by best performance on the validation dataset achieved at any point during the training, where models were evaluated 10 times throughout their training.
To tune our model, we performed a grid search over learning rates in $\{5\times 10^{-3},10^{-3},5\times 10^{-4}\}$ and dropout probabilities in $\{0,0.1,0.2\}$. Overall, we found that learning rates of $10^{-2}$ or larger often caused the model to diverge. We use $6$ transformer layers, $8$ attention heads, and an embedding dimension of $512$ across all settings, except for the Reddit-C and Reddit-S datasets where we use $4$ heads and an embedding dimension of $128$ due to the increased memory cost of these datasets.
@@
Learning Rate Emb. Dim. MLP Dim Heads Transformer Layers Dropout
Hawkes1 $10^{-3}$ $512$ $2048$ 88 66 0.1
Hawkes2 $10^{-3}$ $512$ $2048$ 88 66 0.1
Nonstationary Poisson $10^{-3}$ $512$ $2048$ 88 66 0.1
Nonstationary Renewal $10^{-3}$ $512$ $2048$ 88 66 0.1
Stationary Renewal $10^{-3}$ $512$ $2048$ 88 66 0.1
Self-Correcting $10^{-3}$ $512$ $2048$ 88 66 0.1
PUBG $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
Reddit-C $10^{-3}$ $128$ $256$ 44 66 0.1
Reddit-S $5\times 10^{-3}$ $128$ $256$ 44 66 0.1
Taxi $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
Twitter $10^{-3}$ $512$ $2048$ 88 66 0.1
Yelp-Airport $5\times 10^{-4}$ $512$ $2048$ 88 66 0.1
Yelp-Miss. $10^{-3}$ $512$ $2048$ 88 66 0.1
@@
Learning Rate Emb. Dim. MLP Dim. Heads Transformer Layers Dropout
PUBG $5\times 10^{-3}$ $512$ $2048$ 88 66 $0.1$
Reddit-C $10^{-3}$ $128$ $256$ 44 66 $0.2$
Reddit-S $10^{-3}$ $128$ $256$ 44 66 $0.2$
Taxi $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
Twitter $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
Yelp-Airport $10^{-3}$ $512$ $2048$ 88 66 $0.2$
Yelp-Miss. $5\times 10^{-4}$ $512$ $2048$ 88 66 $0.1$
@@
Learning Rate $\alpha/N_{max}$ Emb. Dim. MLP Dim. Heads Transformer Layers Dropout
PUBG $^{-3} 1.01.0$ $512$ $2048$ 88 66 $0.2$
Reddit-C $10^{-3}$ $1000.0$ $128$ $256$ 44 66 $0.2$
Reddit-S $10^{-3}$ $1000.0$ $128$ $256$ 44 66 $0.0$
Taxi $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
Twitter $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
Yelp-Airport $10^{-3}$ $1000.0$ $512$ $2048$ 88 66 $0.2$
Yelp-Miss. $10^{-3}$ $100.0$ $512$ $2048$ 88 66 $0.2$
@@ Appendix D Additional Details on Baselines
In this section, we provide additional details regarding our baseline methods. All methods are trained at a batch size of $64$ for $1,000$ epochs, using early stopping on the validation set loss. In early experiments, we also evaluated AttNHP [Zuo et al., 2020], a variant of the NHP which uses an attention-based encoder, but found it to be prohibitively expensive in terms of memory cost (requiring more than 24 GB of VRAM) and, as a result, do not include it in our results.
@@
Our first baseline is the intensity-free TPP model of Shchur et al. [2020a]. This model uses an RNN encoder and a mixture of log-normal distributions to parametrize the decoder. We directly use the implementation provided by the authors. [^3] URL: https://github.com/shchur/ifl-tpp. We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and mixture component counts $\{8,16,32,64\}$. Our best hyperparameters can be found in Table 8 and Table 9.
@@
Hawkes1 $10^{-3}$ $10^{-4}$ $32$ 88
Hawkes2 $10^{-2}$ 0 $32$ 88
Nonstationary Poisson $10^{-3}$ $10^{-6}$ $128$ 88
@@
Reddit-C $10^{-2}$ 0 $64$ 88
Reddit-S $10^{-2}$ 0 $64$ $16$
Taxi $10^{-3}$ $10^{-6}$ $128$ 88
Twitter $10^{-2}$ $10^{-5}$ $32$ 88
@@
We additionally compare against the Neural Hawkes Process of Mei and Eisner [2017]. This model uses an LSTM encoder and a parametric form, whose weights are modeled by a neural network, to model the conditional intensity function. In practice, we use the implementation proved by the EasyTPP benchmark [Xue et al., 2024], as this version implements the necessary thinning algorithm for sampling. [^4] URL: https://github.com/ant-research/EasyTemporalPointProcess We perform a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and embedding dimensions in $\{32,64,128\}$. These hyperparameters are chosen as the EasyTPP implementation allows these to be configured easily. Our best hyperparameters are reported in Table 10 and Table 11.
@@
Our diffusion baseline is based on the implementation of Lin et al. [2022], and our decoder model architecture is taken directly from the code of Lin et al. [2022]. [^5] URL: https://github.com/EDAPINENUT/GNTPP At a high level, this model is a discrete-time diffusion model [Ho et al., 2020] trained to generate a single inter-arrival time given a history embedding. Note that as the likelihood is not available in diffusion models, the CDF in the likelihood in Equation (3) is not tractable. Instead, the model is trained by maximizing an ELBO of only the subsequent inter-arrival time.
In preliminary experiments, we found that the codebase provided by Lin et al. [2022] often produced NaN values during sampling, prompting us to make several changes. First, we use the RNN encoder from Shchur et al. [2020a], i.e. the same encoder as the IFTPP baseline, to reduce the memory requirements of the model. Second, we do not log-scale the inter-arrival times as suggested by Lin et al. [2022], as we found that this often led to overflow and underflow issues at sampling time. Third, we do not normalize the data via standardization (i.e., subtracting off the mean inter-arrival time and dividing by the standard deviation), but rather, we scale the inter-arrival times so that they are in the bounded range [1,1][-1,1]. This is aligned with standard diffusion implementations [Ho et al., 2020], and allows us to perform clipping at sampling time to avoid the accumulation of errors. With these changes, our diffusion baseline is competitive, and able to obtain stronger results than previous work has reported [Lüdke et al., 2023].
We use $1000$ diffusion steps and the cosine beta schedule [Nichol and Dhariwal, 2021], and we train the model on the simplified $\epsilon$-prediction loss of Ho et al. [2020]. We train for $1,000$ epochs with early stopping based on the validation set loss. To tune this baseline, we performed a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$, weight decays in $\{0,10^{-6},10^{-5},10^{-4}\}$, history embedding dimensions $\{32,64,128\}$, and layer numbers $\{2,4,6\}$. Our best hyperparameters can be found in Table 12 and Table 13.
@@
Hawkes1 $10^{-3}$ $10^{-6}$ $64$ 22
Hawkes2 $10^{-2}$ $10^{-5}$ $64$ 44
Nonstationary Poisson $10^{-3}$ $10^{-5}$ $128$ 22
Nonstationary Renewal $10^{-3}$ $10^{-4}$ $64$ 22
Stationary Renewal $10^{-2}$ 0 $32$ 66
Self-Correcting $10^{-3}$ 0 $32$ 66
PUBG $10^{-3}$ 0 $64$ 22
Reddit-C $10^{-3}$ $10^{-6}$ $128$ 44
Reddit-S $10^{-3}$ 0 $64$ 44
Taxi $10^{-2}$ 0 $128$ 44
Twitter $10^{-3}$ $10^{-4}$ $64$ 66
Yelp-Airport $10^{-2}$ 0 $32$ 22
Yelp-Miss. $10^{-2}$ $10^{-5}$ $128$ 22
@@
PUBG $10^{-4}$ $10^{-5}$ $32$ 66
Reddit-C $10^{-2}$ $10^{-6}$ $64$ 66
Reddit-S $10^{-3}$ 0 $64$ 44
Taxi $10^{-3}$ $10^{-6}$ $32$ 22
Twitter $10^{-4}$ $10^{-5}$ $64$ 66
Yelp-Airport $10^{-4}$ $10^{-5}$ $64$ 66
Yelp-Miss. $10^{-3}$ $10^{-5}$ $32$ 44
@@
We compare to the Add-and-Thin model of Lüdke et al. [2023] as a recently proposed non-autoregressive baseline. We directly run the code provided by the authors without additional modifications. [^6] URL: https://github.com/davecasp/add-thin We do, however, perform a slightly larger hyperparameter sweep than Lüdke et al. [2023], in order to ensure a fair comparison between the methods considered. We train for $1,000$ epochs with early stopping on the validation loss. Tuning is performed via a grid search over learning rates in $\{10^{-4},10^{-3},10^{-2}\}$ and number of mixture components in $\{8,16,32,64\}$. We choose to tune only these hyperparameters in order to follow the implementation provided by the authors. Our best hyperparameters can be found in Table 14 and Table 15.
@@
Nonstationary Renewal $10^{-2}$ 88
Stationary Renewal $10^{-2}$ 88
Self-Correcting $10^{-4}$ 88
PUBG $10^{-3}$ 88